OpenStudy (anonymous):

Solve: 5c − c + 10 = 34 −6 6 11 −11

OpenStudy (anonymous):

5c - c +10 = 34 transpose 10 5c - c = 34 - 10 4c = 24 divide the whole equation by 4 c=6

OpenStudy (anonymous):

;)